# LibEER: A Comprehensive Benchmark and Algorithm Library for EEG-based Emotion Recognition

### v2

@@ Abstract
EEG-based emotion recognition (EER) has gained significant attention due to its potential for understanding and analyzing human emotions. While recent advancements in deep learning techniques have substantially improved EER, the field lacks a convincing benchmark and comprehensive open-source libraries. This absence complicates fair comparisons between models and creates reproducibility challenges for practitioners, which collectively hinder progress. To address these issues, we introduce LibEER, a comprehensive benchmark and algorithm library designed to facilitate fair comparisons in EER. LibEER carefully selects popular and powerful baselines, harmonizes key implementation details across methods, and provides a standardized codebase in PyTorch. By offering a consistent evaluation framework with standardized experimental settings, LibEER enables unbiased assessments of over ten representative deep learning models for EER across the four most widely used datasets. Additionally, we conduct a thorough, reproducible comparison of model performance and efficiency, providing valuable insights to guide researchers in the selection and design of EER models. Moreover, we make observations and in-depth analysis on the experiment results and identify current challenges in this community. We hope that our work will not only lower entry barriers for newcomers to EEG-based emotion recognition but also contribute to the standardization of research in this domain, fostering steady development. The library and source code are publicly available at https://github.com/XJTU-EEG/LibEER.
@@ I Introduction
Emotion represents a spectrum of subjective cognitive experiences that forms the core framework for interpersonal relationships and serves as a foundation for evaluating human behaviors [1]. Due to its non-invasive nature, high temporal resolution, portability, and cost-effectiveness, electroencephalography (EEG) has become a prominent physiological signal in emotion analysis [30, 31]. As a result, EEG-based emotion recognition (EER) has attracted significant research interest in recent years, with applications spanning diverse fields such as healthcare [32, 41], advertising [33], and education [34].
Traditional approaches to EER primarily rely on hand-crafted features, such as power spectral density (PSD)[40] and differential entropy (DE)[16]. Recently, with advancements in automatic feature engineering, deep learning models have emerged as a promising research area in EER and achieved dominant performance [17, 19, 2]. These models generally target to extract temporal features within each EEG channel, spatial features across different EEG channels, and spectral features in the frequency domain.
@@
  • I3 (Lack of in-depth analysis of SOTA EER research): While deep learning has dominated EER research, many survey papers  have been published summarizing progress from various perspectives. However, these surveys typically report findings without hands-on experience in reproducing results, which may lead to misdirected future research. This lack of practical insight can obscure real challenges faced in the field. Therefore, an in-depth analysis of SOTA EER research based on reproduction results is essential for uncovering the true challenges and advancing the field.
In response to these challenges, we propose LibEER (Library for EEG-based Emotion Recognition) and conduct an analysis of SOTA EER research based on reproduction results with LibEER. To establish a comprehensive benchmark in EER (addressing issue I1), we standardize datasets, evaluation metrics, and experimental settings. Specifically, we review recent SOTA EER studies to identify the most commonly used datasets, SEED [23], SEED-IV [25], DEAP [22], and MAHNOB-HCI [24], and evaluate the corresponding metrics. Our benchmark includes both subject-dependent and subject-independent settings, along with a variety of preprocessing methods. In contrast to the simplified train/test split commonly used in the literature, we adopt a more rigorous train/validation/test split to ensure a more reliable performance evaluation. To address issue I2, we develop the LibEER by systematically selecting and implementing ten representative deep learning-based EER methods, all within a unified training pipeline in PyTorch. LibEER offers a wide range of EEG signal preprocessing functions, evaluation protocols, and experimental scenarios. To enhance its flexibility, we provide user-friendly interfaces that allow researchers to easily customize experimental configurations. Users can set hyperparameters, choose preprocessing methods, and modify backbone models to suit their specific needs. Based on the reproduction results from LibEER, we conduct a series of comprehensive comparison experiments and identify four key observations (addressing issue I3). Through these experiments, we highlight both the real limitations and the progress made in current studies.
It is worth noting that recent research [3] has made important contributions by providing a toolbox for EER. However, our work differs significantly and builds upon these contributions in several key areas. First, we systematically analyze the challenges associated with making fair comparisons in EER tasks, a discussion notably absent in [3]. Second, while the existing benchmark [3] includes a variety of methods, many of them are not specifically tailored for EER tasks. In contrast, we have selected the ten most representative deep learning-based EER methods, based on their impact in this field, making our benchmark better reflect the true SOTA. Finally, the existing benchmark [3] predominantly focuses on research that has already made its code publicly available. However, many influential EER studies do not provide open-source code, which has led to the exclusion of some significant models in their work. In response, we have carefully reproduced all representative studies, including those without publicly available code and rigorously validated the results to ensure their accuracy. We believe that our benchmark and library offer fresh perspectives that will significantly advance the field.
@@
  • We have pioneered the establishment of the first benchmark specifically designed to enable fair comparisons in EEG-based emotion recognition (EER). This benchmark not only reveals the true progress of the EER task but also provides a convenient reference for conducting comparative experiments.
  • We have introduced LibEER as an open-source project on GitHub, offering flexible customization of experimental schemes through user-friendly interfaces. LibEER supports fair and comprehensive evaluation of over ten popular models across four datasets. Serving both as a toolbox and a platform, it assists practitioners in efficiently utilizing and reproducing EER methods.
@@
The paper is organized as follows. Section II provides the background of the EER task. Section III presents the design of the benchmark. Section IV details the library. Section V covers experimental results and in-depth analysis. Section VI identifies future work, and Section VII concludes the whole work.
@@ II-A Data Collection Protocol
The general protocol for EER data collection is illustrated in Fig. 1. Each data collection begins with a hint of start shown to the participant. Afterwards, various stimuli, such as images, audio, or video, are presented to evoke specific emotions in the participant, such as happiness, sadness, or fear. After each stimulus, the participant is asked to self-assess their emotional state, providing a subjective emotion score based on their experience during the stimulus. This process is repeated a set number of times with different stimuli to collect enough data.
In this context, the participant exposed to the stimuli, from whom data is subsequently collected, is commonly referred to as a subject. A continuous segment of EEG data collected from a subject under a specific stimulus condition is called a trial, typically lasting between one and five minutes. Between trials, the subject takes a rest. A series of trials is termed a session. Since EEG data from the same subject, even in the same emotional state, may vary significantly over time, datasets generally include several sessions, with an interval of more than 24 hours between consecutive ones. Given the high cost of EEG data acquisition, a trial is often divided into segments. Two consecutive segments within a trial can be overlapping or non-overlapping. Each segment is called a sample, which serves as the fundamental unit for EER. Notably, within a single trial, all samples should be designated for either training or test. A trial-dependent setting, where some samples within a trial are used for training and others for testing, is impractical for real-world applications and should be considered an invalid setup.
@@ II-B Problem Formulation
Datasets SEED / SEED-IV / DEAP / MAHNOB-HCI
@@
Figure 2: The distinct utilization of EEG signals by different methods
@@
Model Baseline frequency Citation count Publication Publish year Type
@@
CDCN[9] 2 142 TCDS 2020 CNN
EEGNet[8] 7 2979 JNE 2018 CNN
@@
BiDANN[36] 4 250 TAFFC 2021 RNN
ACRNN[11] 3 338 TAFFC 2020 RNN
MS-mda[6] 6 95 Front. Neurosci. 2021 DNN
@@
RGNN[13] 11 545 TAFFC 2020 GNN
GCBNet[15] 3 302 TAFFC 2019 GNN
DGCNN[12] 18 1105 TAFFC 2018 GNN
HSLT[7] 2 79 IEEE SENS J 2022 Transformer
@@ III-A Baseline Selection
  • These 20 methods, based on model architecture, are divided into five categories: CNN, RNN, GNN, DNN, and Transformer. Finally, we selected models with replication results within a 10% error margin compared to the original reported results, while ensuring that each category includes at least one representative model.
Based on these steps, we ultimately identified the twelve most representative works [8, 9, 10, 11, 12, 13, 15, 5, 6, 7, 36, 14] to include in LibEER. TABLE II presents detailed information on these twelve methods. In future work, we will continue to maintain and expand this collection by incorporating more recent and diverse methods. The following subsections will offer detailed descriptions of the five categories, as well as an examination of each selected method.
@@ III-A1 CNN-based Methods
Convolutional Neural Networks (CNNs) are a type of feedforward neural network characterized by their convolutional operations. When applied to EEG signals, CNNs usually treat EEG data as a two-dimensional (2D) representation: one dimension corresponds to the channels, while the other dimension represents either the temporal or spectral features of the EEG data. This structure is illustrated in Fig. 2 (a), which could handle temporal/spectral and spatial information simultaneously. On the other hand, some work relies on 1D CNN to extract temporal and spatial features separately.
LibEER includes three representative CNN-based methods: EEGNet [8], CDCN [9], and TSception [10]. Although not initially proposed for EER, EEGNet is included as a representative 2D CNN-based model due to its significant influence, with over 3000 citations. Both CDCN and TSception employ 1D CNNs for spatial and temporal feature extraction. TSception, a more advanced approach, incorporates multi-scale 1D CNNs for temporal features and employs hemisphere kernels for spatial features.
@@ III-A2 RNN-based Methods
Recurrent Neural Networks (RNNs) are inherently designed for sequential data, making them well-suited for extracting temporal features from EEG signals. For spatial feature extraction, RNNs treat channels as sequence nodes, as illustrated in Fig. 2 (b). In practice, Long Short-Term Memory (LSTM) is a widely used variant of RNNs.
Our benchmark includes three RNN-based methods: ACRNN [11], BiDANN [36], and R2G-STNN [14]. Both ACRNN and BiDANN use LSTM for temporal feature extraction, while R2G-STNN applies bidirectional LSTM to capture spatial features. An attention mechanism is often integrated with LSTM to enhance feature extraction; for instance, ACRNN uses channel-wise attention for spatial features and extended self-attention with LSTM for temporal feature extraction.
@@ III-A3 GNN-based Methods
Graph Neural Network (GNN) is a type of neural network designed to handle data structured as graphs. A graph consists of nodes and edges, which can represent complex relationships and structures. When processing EEG data, the EEG channels are typically treated as nodes in a graph, as shown in Fig. 2 (c). Since EEG data does not inherently contain information about the edges or the adjacency matrix, using graph neural networks to process EEG data often involves leveraging prior knowledge about the relationships between channels or using learnable adjacency matrices as the graph’s adjacency matrix.
We introduced three GNN-based methods. Dynamical Graph CNN (DGCNN)  [12] models multichannel EEG features using a dynamic adjacency matrix that adapts during training. It employs Chebyshev graph convolution for efficient filtering calculations. Regularized GNN (RGNN)  [13] utilizes a neuroscience-inspired adjacency matrix and introduces simple graph convolution networks (SGC). RGNN enhances robustness through node-wise domain adversarial training and emotion-aware distribution learning to address cross-subject variations and noisy labels. Graph Convolutional Broad Network (GCBNet)  [15] combines graph convolutional layers with traditional convolutional layers for effective feature extraction. GCBNet uses Chebyshev graph convolution followed by stacked convolutional layers to abstract high-level features. The GCBNet+BLS variant incorporates the Broad Learning System to enhance feature extraction and classification further.
@@ III-A4 DNN-based Methods
Deep Neural Network (DNN) is a feedforward artificial network consisting of fully connected neurons with a nonlinear activation function, including the input layer, multiple hidden layers, and the output layer. In the field of EER, DNN processes EEG data by transforming and mapping it into the feature space, enabling the extraction and utilization of complex patterns within the data, illustrated in Fig. 2 (d).
In this paper, we have implemented two popular DNN-based methods. Deep Belief Networks (DBN) [5] are probabilistic generative models representing input data distributions through hidden variables. They consist of stacked restricted Boltzmann machines (RBMs), with the output of one RBM serving as input for the next, and use an efficient greedy layer-wise algorithm for pre-training. Multi-Source Marginal Distribution Adaptation (MS-MDA) [6] utilizes MLPs for feature extraction and classification while employing domain adaptation to reduce variance across subjects and sessions. MS-MDA maintains independent branches rather than merging source domains, aligning each with the target domain individually to preserve marginal distributions and enhance adaptation performance.
III-A5 Transform-based Methods
Transformer uses a self-attention mechanism to capture dependencies between elements in a sequence, effectively handling long-range dependencies. As shown in Fig. 2 (e), the EEG data are typically segmented into multiple patches on channels, regions, or temporal segments when using transformers to process EEG data. These patches are then fed into the transformer model for further analysis.
The Hierarchical Spatial Learning Transformer (HSLT) [7] is a hierarchical transformer that extracts discriminative features by capturing EEG spatial dependencies from electrode to brain-region level. It inputs electrode-extracted patches into the first layer of the transformer encoder to aggregate information within the same brain region. The output then serves as input for the second layer, which captures spatial dependencies between different brain regions and highlights those with greater contributions.
@@ III-B2 Splitting
To mitigate these shortcomings, we propose a systematic and rational data-splitting methodology in our benchmark. Specifically, for tasks that are subject-dependent, we allocate each individual’s data into training, validation, and test sets in a 0.6:0.2:0.2 ratio, adjusting as necessary for datasets that do not divide evenly. In the case of cross-subject tasks, we similarly partition subjects into training, validation, and test sets following the same 0.6:0.2:0.2 ratio. Furthermore, our benchmark rigorously adheres to the cross-trial principle, ensuring that samples from the same trial do not appear in both the training and test sets.
@@ III-C Experimental Tasks
The field of EER encompasses various experimental tasks, each designed to address distinct challenges with corresponding models. The primary experimental tasks are categorized as follows: (1) Subject-dependent: This task evaluates model performance for individual subjects, necessitating both training and testing data from the same subject. (2) Cross-subject: This task aims to generalize model performance across different subjects, requiring training and testing data from different subjects. (3) Subject-independent: This task assesses model performance across a group of subjects, without needing training and testing data from the same or different subjects. (4) Cross-session: This task examines model performance across different sessions of the same subject, requiring training and testing data from different sessions of the same individual.
The subject-independent task lies between the subject-dependent and cross-subject tasks, with the latter two being more aligned with real-world applications. The cross-session task necessitates datasets with multiple sessions, which many datasets lack. Therefore, our benchmark focuses on subject-dependent and cross-subject tasks, as these are the most commonly utilized and hold significant practical relevance. Additionally, it is important to note that we treat data from different sessions of the same subject as if they originate from distinct subjects. We provide specific definitions for the subject-dependent and cross-subject tasks below.
@@
Figure 3: The framework and data pipeline of LibEER

### v3

@@ Abstract
EEG-based emotion recognition (EER) has gained significant attention due to its potential for understanding and analyzing human emotions. While recent advancements in deep learning techniques have substantially improved EER, the field lacks a convincing benchmark and comprehensive open-source libraries. This absence complicates fair comparisons between models and creates reproducibility challenges for practitioners, which collectively hinder progress. To address these issues, we introduce LibEER, a comprehensive benchmark and algorithm library designed to facilitate fair comparisons in EER. LibEER carefully selects popular and powerful baselines, harmonizes key implementation details across methods, and provides a standardized codebase in PyTorch. By offering a consistent evaluation framework with standardized experimental settings, LibEER enables unbiased assessments of seventeen representative deep learning models for EER across the six most widely used datasets. Additionally, we conduct a thorough, reproducible comparison of model performance and efficiency, providing valuable insights to guide researchers in the selection and design of EER models. Moreover, we make observations and in-depth analysis on the experiment results and identify current challenges in this community. We hope that our work will not only lower entry barriers for newcomers to EEG-based emotion recognition but also contribute to the standardization of research in this domain, fostering steady development. The library and source code are publicly available at https://github.com/XJTU-EEG/LibEER.
@@ I Introduction
Emotion represents a spectrum of subjective cognitive experiences that underpins interpersonal relationships and human behavior assessment [1]. Electroencephalography (EEG), owing to its non-invasive nature, high temporal resolution, portability, and cost-effectiveness, has become a widely used physiological signal in emotion analysis [37, 38]. Consequently, EEG-based emotion recognition (EER) has garnered increasing attention across domains such as healthcare [39, 48], advertising [40], and education [41]. Early EER methods focused on hand-crafted features like power spectral density (PSD)[47] and differential entropy (DE)[21], while recent advances in deep learning have enabled automatic extraction of temporal, spatial, and spectral features, leading to state-of-the-art performance [22, 24, 2].
@@
  • I3 (Lack of in-depth analysis of SOTA EER research): While deep learning has dominated EER research, many survey papers [2, 4] have been published summarizing progress from various perspectives. However, these surveys typically report findings without hands-on experience in reproducing results, which may lead to misdirected future research. This lack of practical insight can obscure real challenges faced in the field. Therefore, an in-depth analysis of SOTA EER research based on reproduction results is essential for uncovering the true challenges and advancing the field.
To address these issues, we propose LibEER (Library for EEG-based Emotion Recognition), an open-source platform that enables reproduction and analysis of SOTA EER studies. To establish a benchmark (addressing I1), we standardize datasets, evaluation metrics, and experimental settings. Based on a review of recent studies, we include six widely used datasets, SEED [28], SEED-IV [30], DEAP [27], MAHNOB-HCI [29], SEED-V [31], and MPED [32], and support both subject-dependent and cross-subject configurations, using a rigorous train/validation/test split. To address I2, we implement seventeen representative deep learning-based EER methods within a unified PyTorch training pipeline. LibEER provides diverse EEG preprocessing tools, evaluation protocols, and flexible experimental configurations via user-friendly interfaces. Finally, to address I3, we conduct extensive reproduction-based comparisons using LibEER and derive four key observations that reveal the strengths and limitations of existing models. While recent efforts like [3] have introduced EER toolkits, our work differs by providing more in-depth benchmarking analysis, explicitly focusing on EER-specific models, and systematically reproducing several influential methods that lack open-source implementations. These contributions together offer a more comprehensive and practical foundation for advancing EER research.
@@
  • We have pioneered the establishment of the first benchmark specifically designed to enable fair comparisons in EEG-based emotion recognition (EER). This benchmark not only offers an objective reflection of the actual progress in the EER field but also serves as a practical reference for conducting standardized comparative experiments.
  • We have introduced LibEER as an open-source project on GitHub, offering flexible customization of experimental schemes through user-friendly interfaces. LibEER supports fair and comprehensive evaluation of seventeen popular models across six datasets. Serving both as a toolbox and a platform, it assists practitioners in efficiently utilizing and reproducing EER methods.
@@
The paper is organized as follows. Section II provides the background and formal definition of the EER task. Section III presents the design of the benchmark. Section IV details the library. Section V covers experimental results and in-depth analysis. Section VI identifies future work, and Section VII concludes the whole work.
@@ II-A Data Collection Protocol
The general protocol for EER data collection is illustrated in Fig. 1. Each data collection begins with a start hint shown to the participant. Afterwards, various stimuli, such as images, audio, or video, are presented to evoke specific emotions in the participant, such as happiness, sadness, or fear. After each stimulus, the participant is asked to self-assess their emotional state, providing a subjective emotion score based on their experience during the stimulus. This process is repeated a set number of times with different stimuli to collect enough data.
In this context, the participant exposed to the stimuli, from whom data is subsequently collected, is commonly referred to as a subject. A continuous segment of EEG data collected from a subject under a specific stimulus condition is called a trial, typically lasting between one and five minutes. Between trials, the subject takes a rest. A series of trials is termed a session. Since EEG data from the same subject, even in the same emotional state, may vary significantly over time, datasets generally include several sessions, with an interval of more than 24 hours between consecutive ones. Given the high cost of EEG data acquisition, a trial is often divided into segments. Two consecutive segments within a trial can be overlapping or non-overlapping. Each segment is called a sample, serving as the fundamental unit for EER. Notably, within a single trial, all samples should be designated for either training or test. Splitting samples from the same trial across training and test sets is impractical for real-world applications and considered invalid.
@@ II-B Problem Formulation
Datasets SEED / SEED-IV / DEAP / MAHNOB-HCI / SEED-V / MPED
@@
Figure 2: The distinct utilization of EEG signals by different methods.
@@
Model Base. freq. Citation count Publication Publish year Type
EEGNet[11] 7 2979 JNE 2018 CNN
CDCN[12] 2 142 TCDS 2020 CNN
@@
ACRNN[14] 3 338 TAFFC 2020 RNN
BiDANN[43] 4 250 TAFFC 2021 RNN
@@
DGCNN[15] 18 1105 TAFFC 2018 GNN
GCBNet[18] 3 302 TAFFC 2019 GNN
GCBNet_BLS[18] 3 302 TAFFC 2019 GNN
RGNN[16] 11 545 TAFFC 2020 GNN
NSAL-DGAT [20] - 1 TAFFC 2025 GNN
@@
DAN [8] - 199 ICONIP 2018 DNN
DANN [8] - 199 ICONIP 2018 DNN
MS-MDA[7] 6 95 Front. Neuros. 2021 DNN
PR-PL [9] - 55 TAFFC 2024 DNN
HSLT[10] 2 79 IEEE SENS J 2022 Trans.
@@ III-A Baseline Selection
  • These 20 methods, following the common practice in previous studies of categorizing deep learning models by their main architecture [66, 67], are divided into five categories: CNN, RNN, GNN, DNN, and Transformer. Finally, we selected models with replication results within a 10% error margin compared to the original reported results, ensuring that each category includes at least one representative model.
Based on these steps, we ultimately identified the twelve most representative works [11, 12, 13, 14, 15, 16, 18, 6, 7, 10, 43, 17] to include in LibEER. In addition, we include two classic transfer learning methods (DAN and DANN [8]) that have demonstrated superior performance in EER, as well as two recent state-of-the-art approaches (NSAL-DGAT [20] and PR-PL [9]). These methods are also categorized into the corresponding architecture-based groups according to their backbone model structures. TABLE II presents detailed information on these seventeen methods. In future work, we will continue to maintain and expand this collection by incorporating more recent and diverse methods. The following subsections will offer detailed descriptions of the five categories, as well as an examination of each selected method.
@@ III-A1 CNN-based Methods
Convolutional Neural Networks (CNNs) are a type of feedforward neural network characterized by their convolutional operations. When applied to EEG signals, CNNs usually treat EEG data as a two-dimensional (2D) representation: one dimension corresponds to the channels, while the other dimension represents either the temporal or spectral features of the EEG data. This structure is illustrated in Fig. 2 (a), which could handle temporal/spectral and spatial information simultaneously. On the other hand, some work relies on 1D CNN to extract temporal and spatial features separately. LibEER includes three representative CNN-based methods: EEGNet [11], CDCN [12], and TSception [13].
@@ III-A2 RNN-based Methods
Recurrent Neural Networks (RNNs) are inherently designed for sequential data, making them well-suited for extracting temporal features from EEG signals. For spatial feature extraction, RNNs treat channels as sequence nodes, as illustrated in Fig. 2 (b). In practice, Long Short-Term Memory (LSTM) is a widely used variant of RNNs. Our benchmark includes three RNN-based methods: ACRNN [14], BiDANN [43], and R2G-STNN [17].
@@ III-A3 GNN-based Methods
Graph Neural Network (GNN) is a type of neural network designed to handle data structured as graphs. A graph consists of nodes and edges, which can represent complex relationships and structures. When processing EEG data, the EEG channels are typically treated as nodes in a graph, as shown in Fig. 2 (c). Since EEG data does not inherently contain information about the edges or the adjacency matrix, using graph neural networks to process EEG data often involves leveraging prior knowledge about the relationships between channels or using learnable adjacency matrices as the graph’s adjacency matrix. Our benchmark includes five GNN-based methods: DGCNN [15], GCBNet [18], GCBNet_BLS [18], RGNN [16], and NSAL-DGAT [20].
@@ III-A4 DNN-based Methods
Deep Neural Network (DNN), in this paper, specifically refers to the feedforward fully connected neural network, also known as Fully Connected Neural Network (FNN) or Multi-Layer Perceptron (MLP). It consists of an input layer, multiple hidden layers composed of fully connected neurons with nonlinear activation functions, and an output layer. Unlike other deep learning architectures such as CNN, RNN, or GNN, DNN focuses on dense layer transformations without convolutional, recurrent, or graph-based structures. In the field of EER, DNN processes EEG data by performing feature transformation and mapping into a high-dimensional feature space, enabling the extraction and utilization of complex patterns within the data, as illustrated in Fig. 2 (d). Our benchmark includes five DNN-based methods: DBN [6], DAN [8], DANN [8], MS-MDA [7], and PR-PL [9].
III-A5 Transformer-based Methods
Transformer uses a self-attention mechanism to capture dependencies between elements in a sequence, effectively handling long-range dependencies. As shown in Fig. 2 (e), the EEG data are typically segmented into multiple patches on channels, regions, or temporal segments when using transformers to process EEG data. These patches are then fed into the transformer model for further analysis. Our benchmark includes one Transformer-based methods: HSLT [10].
@@ III-B2 Splitting
To mitigate these shortcomings, we propose a systematic and rational data-splitting methodology in our benchmark. Specifically, for tasks that are subject-dependent, we allocate each individual’s data into training, validation, and test sets in a 0.6:0.2:0.2 ratio, adjusting as necessary for datasets that do not divide evenly. The SEED-V dataset contains relatively fewer trials and more classes, so the ratio was adjusted to 1:1:1. In the case of cross-subject tasks, we similarly partition subjects into training, validation, and test sets following the same 0.6:0.2:0.2 ratio. Furthermore, our benchmark rigorously adheres to the cross-trial principle, ensuring that samples from the same trial do not appear in both the training and test sets.
@@ III-C Experimental Tasks
The field of EER encompasses various experimental tasks, each designed to address distinct challenges with corresponding models. The primary experimental tasks are categorized as follows: (1) Subject-dependent: This task evaluates model performance for individual subjects, requiring both training and testing data from the same subject. (2) Cross-subject: This task aims to generalize model performance across different subjects, requiring training and testing data from different subjects. (3) Subject-independent: This task assesses model performance without constraining the subject-level split, allowing training and testing data to be randomly divided from mixed samples of multiple subjects. (4) Cross-session: This task examines model performance across different sessions of the same subject, requiring training and testing data from different sessions of the same individual. Figure 3 provides an intuitive illustration of the differences among the four tasks.
Figure 3: Four different experimental tasks in EER.
In previous studies, the definitions of subject-independent and cross-subject tasks have often overlapped, as subject-independent is typically implemented using leave-one-subject-out cross-validation (LOSO-CV) [15, 16]. To avoid ambiguity, we explicitly adopt cross-subject as one of the core evaluation settings in our benchmark. Cross-subject strictly requires training and testing on entirely different subjects, emphasizing the model’s generalization ability. This task setup is more aligned with real-world applications where models are expected to perform reliably on new, unseen users in large-scale deployment scenarios. In addition, subject-dependent tasks are included to evaluate a model’s performance under personalized settings, where training and testing are performed on the same subject. This setup is important for scenarios focusing on optimizing accuracy for specific individuals, such as personalized healthcare or user-specific emotion monitoring. Therefore, our benchmark concentrates on cross-subject and subject-dependent tasks as the most widely adopted and practically meaningful evaluation setups. Additionally, it is important to note that we treat data from different sessions of the same subject as if they originate from distinct subjects. We provide specific definitions for the subject-dependent and cross-subject tasks below.
@@
SEED-V [31] 3 16 15 62 movie clip 120s-240s 1000 raw data happy/sad/disgust/netural/fear
MPED [32] 1 23 28 62 movie clip 150s-300s 1000 STFT joy/funny/anger/fear/disgust/sad/neutrality